OpenStudy (anonymous):

use implicit differentiation to find dy/dx at the point (1,1) if sqrt(xy)=1+x^2y

OpenStudy (experimentx):

(y+xdy/dx)/sqrt(xy) = 2xy+x^2dy/dx 1+dy/dx = 2 + dy/dx ??

OpenStudy (experimentx):

1/2*(y+xdy/dx)/sqrt(xy) = 2xy+x^2dy/dx 1/2 - 2 = 1/2 * dy/dx

OpenStudy (anonymous):